OpenStudy (anonymous):

Evaluate the discriminant of 25x2 + 350x + 49 = 0. Tell how many solutions the equation has and whether the solutions are real or imaginary. A. one real solution B. two imaginary solutions C. two real solutions D. cannot be determined

OpenStudy (anonymous):

\[ 25x^2 + 350x + 49 = 0\] \[ax^2+bx+c=0\] discriminant is \(b^2-4ac\) in your case it is \[305^2-4\times 25\times 49\]whatever that is

OpenStudy (anonymous):

if it is positive, two solutions negative, no real solution zero, one solution namely \(\frac{b}{2a}\)

OpenStudy (anonymous):

isn't it \[-\frac{ b }{ 2a }\] if the discriminant is zero or did I miss something?